OpenStudy (anonymous):

Why is anything to the 0 power =1

OpenStudy (anonymous):

\[a^0=a^{k-k}=a^k*a^{-k}=\frac{a^k}{a^k}=1\]

OpenStudy (anonymous):

for any number except 0